OpenStudy (anonymous):

Two resistors are connected in parallel across an ideal 12.0 Volts battery. Resistor A has a value of 20 ohm and resistor B carries a current of 0.2810 A. What is the potential difference across resistor A? I have been trying to use V=IR but idk if that is the correct formula

OpenStudy (anonymous):

yes, find R(B) using R=V/i, v=12 volts and i= 0.2810

OpenStudy (anonymous):

I did 12/.2810 and got 42.704

OpenStudy (anonymous):

if your only looking for the potentioal difference across A ,,since they are parallel plates than the potential difference is the same... so the answer would be 12 volts

OpenStudy (anonymous):

theres like 4 other questions that go along with it I just am confused here are the other ones: What is the potential difference across resistor B? What is the current in resistor A? What is the resistance of B? What is the total power dissipated in the circuit?

OpenStudy (anonymous):

ok the potential difference is the same for A and B which is 12 volts,,,current in resistor A=V/R,12/20 = 0.6 Amp. resistance B you found above V/i,,12/0.2810 = 42.704 ohms...total power dissipated would be sum of V/R(A) + V/R(B)...12/20 plus 12/42.704 and the units are in W (watts)

OpenStudy (anonymous):

oh ohms! i was using volts an it was incorrect lol